Evaluate the integral. dt 11 (1+) 5:2 (1+)? dt =D
Answer #1

Evalute the integral = ( 47 (1748)t det Soft (1748) de ☆ 5 (v++8) + d(484) too 7 = = + + c 8 7+1 8 a Ś (1-1)* te 8 = (ITA) +C